# The Fate of Oceans on First-Generation Planets Orbiting White Dwarfs

### v1

@@ 2.2 A Model for Ocean Evaporation due to Irradiation
[…]
@@ 3.2 Thermal Effects of the Circularizing Orbit
$\begin{split}\frac{da}{dt}=-\sqrt{\frac{GM_{\star}^{3}}{a^{11}}}\,\bigg{[}\frac{9}{2}\frac{\zeta}{Q_{\star}}M_{\star}^{-5/2}R^{5}m_{p}+\\ \frac{21\ k_{2}}{Q_{p}m_{p}}\frac{r_{p}^{5}}{(1-e^{2})^{15/2}}\biggl{(}1+31\frac{e^{2}}{2}+255\frac{e^{4}}{8}+185\frac{e^{6}}{16}\\ -(1-e^{2})^{3/2}\left(1+15\frac{e^{2}}{2}+45\frac{e^{4}}{8}+5\frac{e^{6}}{16}\right)\biggr{)}\bigg{]}\,\end{split}$ (11)
@@
$\begin{split}\frac{de}{dt}=-e\,\sqrt{\frac{GM_{\star}^{3}}{a^{13}}}\,\bigg{[}\frac{171}{16}\frac{\zeta}{Q_{\star}}M_{\star}^{-5/2}R^{5}m_{p}+\\ \frac{21}{2}\frac{k_{2}}{Q_{p}m_{p}}\frac{r_{p}^{5}}{(1-e^{2})^{13/2}}\biggl{(}1+15\frac{e^{2}}{4}+15\frac{e^{4}}{8}+5\frac{e^{6}}{16}\\ -\frac{11(1-e^{2})^{3/2}}{18}\left(1+3\frac{e^{2}}{2}+\frac{e^{4}}{8}\right)\biggr{)}.\end{split}$ (12)
@@
[…]
@@
Assuming that the lithosphere is thin enough that the tidal deformation exactly matches the tidal geopotential, $\overline{\dot{\varepsilon}_{ij}^{2}}$ for planet with eccentricity $e$ is approximately $415\omega^{2}\gamma^{2}e^{2}/$ 4 (Ojakangas & Stevenson, 1989), where $\gamma=(a^{3}\omega^{2})/(Gm_{p})$ characterizes the relative magnitude of tidal forcing and the planet’s self-gravity. Given that the bulk of heat production will take place in the lithosphere, we can obtain the heat flux at the top of the lithosphere $\mathcal{F}_{s}$ from Equation 18 and Equation 16:

### v2

@@ 2.2 A Model for Ocean Evaporation due to Irradiation
$\frac{L_{\text{XUV}}}{L_{\text{bol}}}=\frac{\int_{\text{XUV}}d\nu\>u_{\nu}(\nu,T)}{\int_{0}^{\infty}d\nu\>u_{\nu}(\nu,T)}.$ (4)
@@ 3.2 Thermal Effects of the Circularizing Orbit
$\begin{split}\frac{da}{dt}=-\sqrt{\frac{GM_{\star}^{3}}{a^{11}}}\,\bigg[\frac{9}{2}\frac{\zeta\ k_{2,\star}}{Q_{\star}}M_{\star}^{-2}R^{5}m_{p}+\\ \frac{21\ k_{2,p}}{Q_{p}m_{p}}\frac{r_{p}^{5}}{(1-e^{2})^{15/2}}\biggl(1+31\frac{e^{2}}{2}+255\frac{e^{4}}{8}+185\frac{e^{6}}{16}\\ -(1-e^{2})^{3/2}\left(1+15\frac{e^{2}}{2}+45\frac{e^{4}}{8}+5\frac{e^{6}}{16}\right)\biggr)\bigg]\,\end{split}$ (11)
@@
$\begin{split}\frac{de}{dt}=-e\,\sqrt{\frac{GM_{\star}^{3}}{a^{13}}}\,\bigg[\frac{171}{16}\frac{\zeta\ k_{2,\star}}{Q_{\star}}M_{\star}^{-2}R^{5}m_{p}+\\ \frac{21}{2}\frac{k_{2,p}}{Q_{p}m_{p}}\frac{r_{p}^{5}}{(1-e^{2})^{13/2}}\biggl(1+15\frac{e^{2}}{4}+15\frac{e^{4}}{8}+5\frac{e^{6}}{16}\\ -\frac{11(1-e^{2})^{3/2}}{18}\left(1+3\frac{e^{2}}{2}+\frac{e^{4}}{8}\right)\biggr).\end{split}$ (12)
@@
$\frac{\kappa_{0}}{2}\left.\left(\frac{d\ln T}{dz}\right)^{2}\right|_{\mathrm{bot}}^{\mathrm{surf}}=\int_{T_{s}}^{T_{m}}\frac{q(T)}{T}dT$ (16)
@@
Assuming that the lithosphere is thin enough that the tidal deformation exactly matches the tidal geopotential, $\overline{\dot{\varepsilon}_{ij}^{2}}$ for planet with eccentricity $e$ is approximately $15\omega^{2}\gamma^{2}e^{2}/4$ (Ojakangas & Stevenson, 1989), where $\gamma=(a^{3}\omega^{2})/(Gm_{p})$ characterizes the relative magnitude of tidal forcing and the planet’s self-gravity. Given that the bulk of heat production will take place in the lithosphere, we can obtain the heat flux at the top of the lithosphere $\mathcal{F}_{s}$ from Equation 18 and Equation 16: